Extra Problem 2. Four mutant strains of Neurospora (#1-4) were examined by complementation testing. Results are shown below (+ indicates growth; - indicates no growth).

1

2

3

4

1

-

+

-

+

2

+

-

+

-

3

-

+

-

+

4

+

-

+

-

Crosses among the mutant strains yielded prototrophs in the following percentages:

            1 x 2   25.0%                        2 x 3   25.0%                        3 x 4   25.0%

            1 x 3   0.02%                        2 x 4   none

A. How many genes are mutant among these strains?

B. Give map distances between all pairs of linked mutations and indicate which mutations are unlinked.
Answer #1

Answer A- There are two mutant genes present in these strains. The mutant groups are group1=1,3 (because neither the complementation of 1 or 3 could rescue growth of 1), group2=2,4 (because neither the complementation of 2 or 4 could rescue growth of 4).

Answer B-

The mutations 1 and 3 are linked. Similarly mutations 2 and 4 are also linked

The recombinant frequency of (1-2)=25% (unlinked), (1-3)=~0% (linked), (2-3)=25% (unlinked), (2-4)=0% (linked) and (3-4)=25%(unlinked).

The gene map can be drawn as follows,

Mutations 1 and 2 are unlinked because 1 and 2 assort independently. Similarly mutations 3 and 4 are unlinked because 3 and 4 assort independently.